OpenStudy (anonymous):

Solve the equation on the interval [0, 2π). cos x + 2 cos x sin x = 0

OpenStudy (anonymous):

@Calcmathlete

OpenStudy (cwrw238):

factor the left side cos x (1 + 2 sin x ) = 0

OpenStudy (anonymous):

so its A?

OpenStudy (cwrw238):

so do you know what angles have a cos of 0? angles between o and 2pi?

OpenStudy (cwrw238):

x = pi/2 has a cos of 0

OpenStudy (anonymous):

so is it A?

OpenStudy (cwrw238):

hmmm - you are twisting my arm - 3pi/2 also has a cos of 0 what about the sine bit? 2 sin x = -1 so sin x = -1/2 what angles have a sine of -1/2? clue: sin pi/6 has a sine of 1/2

OpenStudy (cwrw238):

if you are unhappy with radians pi/6 = 30 degrees and sine is negative in 3rd and 4th quadrants

OpenStudy (anonymous):

7pi/6 and 11pi/6 so its B!

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

thnx! srry for twisting ur arm :)

OpenStudy (cwrw238):

lol - i twisted yours in reply

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

i have one more! LAST one i promise :) Solve the equation on the interval [0, 2π). sec(x/2) = cos(x/2) Im not sure how to simplify or factor

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

change sec(x/2) to 1 / cos (x/2) giving 1 / cos(x/2) = cos(x/2)

OpenStudy (anonymous):

ok. then what do i do?

OpenStudy (cwrw238):

so cos (x/2) = ?

OpenStudy (cwrw238):

1/x = x so what is x?

OpenStudy (anonymous):

2 cos (x/2) ?

OpenStudy (cwrw238):

nope

OpenStudy (anonymous):

cos(x/2)^2?

OpenStudy (cwrw238):